# Understanding In-Context Learning with a Pelican Soup Framework

### v1

@@ 2.5 Extending the Framework with Formal Language
There exists (1) a finite set of atom concepts $\Omega$ and (2) a function $f$ that can map any sentence in language to its meaning represented with a logical formula with operands in $\Omega$ such that for any two sentences $s_{1},s_{2}$, the logical relation between $s_{1}$ and $s_{2}$ is the same as $f(s_{1})$ and $f(s_{2})$.
This assumption is aligned with the cognitive psychology studies that hypothesize the existence of a set of mental tokens (Fodor, 1975, 2008) and a recent study (Piantadosi, 2021) showing that semantics can be encoded with the combination of only a few symbols.
This assumption suggests that if we have the parsing function $f$, solving NLP tasks only requires a finite-length program that can do logical reasoning by manipulating logical symbols according to logical induction rules. If a deep model can learn this program, then it can perform a task even if this task is not in the training data. This assumption of a finite $\Omega$ also instantiates the concept of language skills by Arora & Goyal (2023), and their theoretical results are thus applicable.
2.6 Relating to Function Description Length
When there is no decoding constrains, $\Pr(r_{t},d|x_{t},z^{*},S_{t-1})$ is related to the difficulty of the example. To see this, we need an additional assumption:
@@ 2.6 Relating to Function Description Length
In some documents in the training data, the paragraphs are constituted with steps in a logical induction process with some steps randomly dropped.
This kind of document may be pervasive in the training data. Essays arguing some claims are one example. To be convincing, these essays should proceed like a proving process that induces their conclusions. Documents describing a series of events can be another example, as events follow commonsense and develop progressively.
@@
where $\ell(x_{t})$ is the number of reasoning steps required to solve the task, and $c$ is a constant This $\ell(x_{t})$ corresponds to the description length of the function that maps the inputs to their label in the loss bound by Hahn & Goyal (2023) (more discussion in Appendix D).
Train
x57 x56 x64 r3 $\to$ x79 , r1 x57 $\to$ x58 , x90 x58 $\to$ r3 , r3 r1 $\to$ x20 , … ; x80 x66 x63 x83 x1 $\to$ x82 , x80 x82 $\to$ r1 , … , x64 x80 $\to$ x54 .
ICL
x44 x67 x34 x62 $\to$ r2 ; x55 x38 x50 x48 $\to$ r1 ; x21 x59 x57 x86 $\to$ r2 ; x55 x76 x84 x99 $\to$
CoT
x44 x67 x34 x62 $\to$ x16 , x34 x62 $\to$ x99 , x99 x16 $\to$ r1 ; x77 x34 x62 x97 $\to$ x12 … ; x21 x59 x57 x86 $\to$ x69 , x59 x57 $\to$ x75 , x69 x75 $\to$ r2 ; x55 x76 x84 x99 $\to$
Table 1: Calcutec examples for training, in-context learning (ICL), and chain-of-thought (CoT).
3 Distribution Shifts from Language Modeling to In-context Learning
In addition to the theoretical analyses, an important aspect of understanding ICL is to understand how models generalize. The bounds we have established in § 2.4 are based on the analyses of the marginal distribution conditioned on the input $S_{t-1},x_{t}$. Because the distribution of prefix $S_{t-1},x_{t}$ is different from the distribution of general text, a model needs to overcome several types of distribution shifts to predict the marginal distribution precisely. We characterize a few of them below.
@@ 3 Distribution Shifts from Language Modeling to In-context Learning
The demonstration is in the format $x_{t},r_{t},d$, where $x_{t}$ follows the downstream task distribution. However, in the LM training data, sequences in this format are unlikely to happen frequently.
@@
The verbalizers appear very frequently in $S_{t-1}$. However, the verbalizer may not have as high frequency in the training data distribution.
Unseen Reference-Meaning Association.
When performing ICL, the verbalizers are associated with descriptions of the classes for the task. However, the training data may not include this kind of association.
We empirically investigate these types of distribution shifts in the next section.
4 Experimenting with Synthetic Datasets
As a concrete instance of our framework, we present a toy setup, Calcutec, which includes a data generation process for LM training data and a process for downstream tasks. Calcutec follows our framework and mimics the distribution shifts LMs need to overcome as discussed in §3. We will use this setup to inspect models’ generalization ability.
4.1 Calcutec
Setup.
Following the assumption in §2, we construct a pseudo-language with the following components:
  • Logic model: We use a subset of propositional logic as our logic model. We only consider Horn clauses (Horn, 1951), i.e., formulas in the form $A\wedge B\to C$. This allows us to efficiently generate proof with time complexity linear to the size of the knowledge base (Dowling & Gallier, 1984).
  • Atom concepts: We use 100 variables as our set of atom concepts $\Sigma$.
  • $\mathrm{KB}$: For each variable $\sigma\in\Sigma$, there are 5 formulas in the knowledge base whose consequence (the literals after $\to$ in a horn clause) is $\sigma$, while the antecedents (the literals before $\to$) are sampled from $\Sigma\backslash\{\sigma\}$ uniformly at random.
  • We have a set $\Gamma=\{r_{i}\}_{i=1}^{4}$ representing the reference words described in Assumption 2.3.
Training Dataset.
Following Assumption 2.4, a document is a concatenation of paragraphs separated by delimiter “;” and ends with “.”. In our synthetic language model training dataset, each document contains 16 paragraphs. Before starting the generation process, we randomly draw a reference symbol $r_{a}\in\Gamma$ and associate it with two formulas $s_{1},s_{2}$. For simplicity, we make these two formulas single-variable, i.e., $s_{1},s_{2}\in\Sigma$.
Because paragraphs in the real world are sentences ordered in a certain way, we follow Assumption 2.9 and randomly generate a proof as a paragraph. Each paragraph represents the induction process of $P\models g$ for some randomly selected $P\subset\Sigma$ and $g\in\Sigma$ that allows us to use the reference symbol $r$ in the paragraph. Each sentence in the paragraph is a sentence representing a reasoning step. Because we use Horn logic as our logic model, a step is a rule in $\mathrm{KB}$ in the form $a_{1}a_{2}\to b$. We separate the clauses in the sequence with commas. To simulate the fact that documents in the real world always skip some reasoning steps, we further apply some perturbations over the generated paragraphs that drop some reasoning steps with a skip rate $p_{skip}$ (details in Appendix E).  [^2] Models can acquire in-context learning ability even with $p_{skip}=0$ (Figure 6 in thr appendix).
After we generate a document, we replace $s_{1}$ and $s_{2}$ with the reference symbol $r_{a}$. We randomly draw another reference symbol $r_{b}\in\Gamma$ and replace the two most frequent symbols in the document with $r_{b}$. (We provide the pseudo-code Alg. 1 in the appendix.)
Downstream Tasks.
Following the formulation in Assumption 2.5, we define a binary classification task by defining the description $v_{+}$ and $v_{-}$ for the positive and negative classes, respectively. $v_{+}$ and $v_{-}$ are the disjunctions of atom concepts, i.e., in the form of $a_{1}\vee a_{2}\vee\cdots$. We create five downstream tasks using different disjunctions. Each input is a subset of variables in $\Sigma$ from which we ensure that only one of the classes can be induced.
Demonstration.
We represent an input-label pair as $x_{1}x_{2}\cdots\to r$, where $x_{1}x_{2}\cdots$ is the input part and $r\in\{r_{+},r_{-}\}\subset\Gamma$ is a reference symbol serving as a verbalizer.
Chain-of-thought.
A chain-of-thought is in the format same format as the training data, but ends with a reference symbol $r\in\{r_{+},r_{-}\}$, e.g., $x_{1}x_{2}\cdots\to x_{3};x_{3}\cdots x_{4}\to r_{+}$. This chain-of-thought reflects the step-by-step induction process from the inputs to the label.
4.2 Inspecting the Distribution Shifts
Our Calcutec dataset allows us to inspect some types of distribution shifts in § 3.
Format Mismatch.
The training data contains reasoning steps but these steps are absent in the demonstration.
Verbalizer Mismatch.
When we are picking the reference symbols in $\Gamma$, we assign probability mass 45%, 45%, 5%, 5% to $r_{1},r_{2},r_{3},r_{4}$. In this way, we can inspect whether using less frequent verbalizers will lead to worse performance.
@@ 4.2 Inspecting the Distribution Shifts
To investigate whether the model can generalize to a new combination of formulas unseen in the training data when we generate our training data, we ensure that the references are never associated with some combinations of atom concepts ($s_{1},s_{2}\in\Sigma$). We then test the trained model on tasks where $v_{+}$ and $v_{-}$ are the disjunctions of unseen combinations. Additionally, while in the training data, each reference symbol is associated with only two atom concepts; we test whether the model can perform in-context learning on tasks where $v_{+}$ and $v_{-}$ are the disjunctions of three atom concepts.
4.3 Experiment Details
We use $p_{skip}=0.25$ in our experiment. We generate 60,000 documents with 16 paragraphs following the process described in § 4. Among them, we use 10k for validation. We train a 6-layer Transformer (Vaswani et al., 2017) model until the loss on the validation set converges. We include additional setups in §G
4.4 Experiment Results and Discussion
(a) Verbalizers = $r_{1},r_{2}$
(b) Verbalizers = $r_{3},r_{4}$
Figure 1: In-context learning accuracy with Calcutec when using different verbalizers ($r_{1},r_{2}$ or $r_{3},r_{4}$). The dotted lines represent the performance of unseen combinations described in §4.2. The colors represent the number of formulas in $\Sigma$ each class ($v_{+}$ or $v_{-}$) is associated with. The main lines represent the average accuracy of 5 tasks. We plot the performance of each task in lighter colors.
@@ 4.4 Experiment Results and Discussion
Table 2: The 4-shot accuracy of in-context learning (ICL) versus chain-of-thought (CoT) using different verbalizers.
task direct pronoun
SST-2 63.0 65.3
CR 61.7 62.9
MR 59.2 56.7
Subj 51.0 62.2
@@
Training
4 9 2 8 4 6 + 0 8 0 3 5 0 = 0 0 0 0 0 0 ; 0 9 2 8 4 6 + 0 8 0 3 5 0 = 4 0 0 0 0 0 ; 0 0 2 8 4 6 + 0 0 0 3 5 0 = 4 7 1 0 0 0 ; 0 0 0 8 4 6 + 0 0 0 3 5 0 = 4 7 3 0 0 0 ; $\cdots$ 0 0 0 0 0 0 + 0 0 0 0 0 0 = 4 7 3 1 0 7 ;
Testing
8 7 4 0 1 6 + 0 9 2 1 5 0 = 0 0 0 0 0 0 ; 0 0 0 0 0 0 + 0 0 0 0 0 0 =
Table 4: Training and testing examples for the digit addition task.
Figure 1 shows that the model trained with Calcutec can perform in-context learning, indicating that the model can generalize under the distribution shifts described in §3. This evidence supports our Pelican Soup hypothesis. We further inspect the in-context learning performance under the distribution shifts described in §4.2:
  • Infrequent verbalizer: We observe similar performance of using frequent reference symbols ($r_{1},r_{2}$) or infrequent reference symbols ($r_{3},r_{4}$) as the verbalizers.
  • Unseen tasks: Figure 1 shows that the model has similar performance over tasks defined with unseen and unseen combinations of atom concepts (dot lines and solid lines). The models can also generalize to tasks defined with three formulas (green lines).
In sum, the results show that the model can generalize well under several distributional shifts.
We experiment with 4-shot learning using chain-of-thought. Table 2 shows that the model also benefits from chain-of-thought. We conjecture that it is because chain-of-thought has a format more similar to the format of training.
5 Digit Addition Task
In § 4, we show that transformer models can generalize to downstream tasks where the intermediate reasoning steps are missing. As this is related to models’ ability to predict $P(r|x,K)$2.6) when we force the model to skip the intermediate steps, we are curious whether transformer models can still generalize even when more steps are missing. Since in the Calcutec environment, it is difficult to construct a task that requires a long reasoning process without exponentially increasing the size of the atom concept set $\Sigma$, We propose another toy task, the digit addition task, which involves longer reasoning processes.
5.1 Setup
The digit addition task requires the model to perform the addition operation over two $n$-digit numbers. It requires $n$ steps to solve if we solve it digit-by-digit. We generate a training set consisting of digit-by-digit reasoning steps for 100,000 pairs of numbers. Each intermediate step may be dropped at a constant probability $p_{drop}$. After training a Transformer model with the language modeling objective, we test whether the model can generate the final answer without generating the intermediate steps. We experiment with different models of different sizes. From this, we can verify whether models of different sizes can acquire the intuition from the language modeling task.
We show examples of our data in Table 4. Each digit sequence represents a number from the lower digit to the higher digit. The reasoning process in the training set gradually updates both sides of “=” from the lowest digit to the highest digit. As for the testing example, we skip the intermediate steps, making the model complete right-hand side of “=” in the last step. The model needs to acquire intuition from the training set in order to solve the testing samples. (We include a rigorous description in Appendix I.)
Figure 2: The exact accuracy (y-axis, solid points) and digit-level accuracy (y-axis, hollow points) versus validation loss (x-axis) for the 5-digit addition task for five random seeds and different dropping rates $p_{drop}$ (In the appendix, we include the complete results in Figure 11 and the results of the 6-digit addition task in Figure 12.)
5.2 Results and Discussion
We report the exact match accuracy and the digit-level accuracy of models trained with different $p_{drop}$ in Figure 2 with 5 random seeds. A higher accuracy implies the better intuition the model acquires from the sequence modeling task. The results show that three of the four models can achieve perfect accuracy when $p_{drop}$ is as small as $0.3$. It suggests that it is possible that the model can predict the final answer even without intermediate steps, even when examples in the training data are mostly step-by-step. We also observe that larger models tend to have higher and more stable accuracy. When the number of digits is 6 (Figure 12 in the appendix), only the largest model can achieve perfect accuracy. This observation is aligned with the emergence of large language models’ ability.

### v2

@@ 5 Generalization
There exists (1) a finite set of atom concepts $\Sigma$, (2) a knowledge base $\mathrm{KB}$ consisting of logical rules between the atom concepts in $\Sigma$, and (3) a function $f$ that can map any sentence in language to its meaning represented as a logical formula with operands in $\Sigma$ such that for any two sentences $s_{1},s_{2}$, the logical relation between $s_{1}$ and $s_{2}$ judged by humans is the same as $f(s_{1})$ and $f(s_{2})$ given the rules in the knowledge base $\mathrm{KB}$.
The three elements of this assumption correspond to theories in various fields. The notion of atom concepts aligns with cognitive psychology studies that hypothesize the existence of a set of mental tokens Fodor (1975, 2008) and a recent study Piantadosi (2021) suggesting semantics can be encoded with the combination of only a few symbols. The notion of knowledge base follows the early formulation of AI (Siskind, 1996; Murphy, 2004). As for the existence of a parsing function $f$, it follows the long history of linguistics studying the relationships between natural and formal languages (Carnap and others, 1968; Bresnan and Bresnan, 1982; Steedman, 1987, 1996; Sag et al., 1999; Frege and others, 1879; Peirce, 1883).
This assumption suggests that if we have the parsing function $f$, solving NLP tasks only requires a finite-length program that can do logical reasoning by manipulating logical symbols according to logical induction rules. If a model can learn this program, then it can perform a task even if this task is not in the training data. This assumption of a finite $\Sigma$ also instantiates the concept of “language skills” by Arora and Goyal (2023), and their theoretical results are thus applicable.
Figure 1: Calcutec examples for training, in-context learning (ICL), and chain-of-thought (CoT).
6 Relating to Description Length
We can see $\Pr(r_{t},d|x_{t},z^{*},S_{t-1})$ as the difficulty of the example by having an additional assumption:
@@ 6 Relating to Description Length
In some documents in the training data, the paragraphs are constituted with steps in a logical induction process, with some steps randomly dropped.
This kind of document may be prevalent in training data. Essays arguing some claims are one example. To be convincing, these essays should proceed like a proving process that induces their conclusions. Documents describing a series of events can be another example, as events follow commonsense and develop progressively.
@@
where $\ell(x_{t})$ is the number of reasoning steps required to solve the task, and $c$ is a constant. This $\ell(x_{t})$ corresponds to the description length of the function that maps the inputs to their label in the loss bound by Hahn and Goyal (2023) (more discussion in Appendix E).
7 Inspecting Generalization Empirically
Although we have shown that perfect LMs can follow instructions (§3.2) and perform ICL (§4) when training data satisfy the assumptions in §2, it remains unverified whether imperfect real-world LMs exhibit the same capabilities. For LMs to demonstrate these abilities, they must generalize well from the training set to the prompts used during inference. However, analyzing how deep models generalize inherently requires additional assumptions about the model, which is beyond the scope of this paper. Thus, to address this gap, we leverage our extended framework (§5) to characterize distribution shifts and design experiments that empirically examine the generalization of LM.
7.1 Inspecting the ICL Capability
Firstly, we present a synthetic setup, Calcutec, as a concrete instantiation of our theoretical framework. With Calcutec, we show that Transformers can acquire ICL capability by modeling the linguistic characteristics specified in our framework.
7.1.1 Calcutec
Setup
Following the extended framework in §5, we construct a pseudo-language:
  • Logic model: We use a subset of propositional logic as our logic model. We only consider Horn clauses (Horn, 1951), i.e., formulas in the form $A\wedge B\to C$.
  • Atom concepts: We use 100 symbols as our set of atom concepts $\Sigma$.
[…]
  • We have a set $\Gamma=\{r_{i}\}_{i=1}^{4}$ representing the expressions described in Assumption 3.3.
Training Dataset.
Following Assumption 3.4, a document is a concatenation of paragraphs separated by delimiter “;” and ends with “.”. In our synthetic language model training dataset, each document contains 16 paragraphs.
Following Assumption 3.2, each paragraph consists of sentences logically coherent to each other. Because sentences in the real world are not ordered arbitrarily, we follow Assumption 6.1 and generate random paragraphs following the structure of logical proofs. Each paragraph represents an induction process of $P\models g$ for some randomly selected $P\subset\Sigma$ and $g\in\Sigma$. Each sentence in the paragraph is a sentence representing a reasoning step for $P\models g$. We separate the sentences in the sequence by commas. To simulate the noise in the real world, we further apply perturbations that skip some steps with a skip rate $p_{skip}$.  [^1] Models can acquire in-context learning ability even with $p_{skip}=0$ (Figure 6 in the appendix).
Following Assumption 3.3, after we generate a document, we randomly associate some symbols $A,B\subset\Sigma$ with $r_{a},r_{b}\in\Gamma$ respectively. We reflect this association in the generated document by replacing symbols in AA and $B$ with expression $r_{a},r_{b}\in\Gamma$ respectively (details in Appendix F.1).
Downstream Tasks.
Following the formalism in §3.2, we define a binary classification task by defining the descriptions $v_{+}$ and $v_{-}$ of the positive and negative classes, respectively. We use the disjunctions of atom concepts (i.e. in the form of $a_{1}\vee a_{2}\vee\cdots$) as descriptions of classes. We create five downstream tasks using different disjunctions. Each input is a subset of variables in $\Sigma$ from which we ensure only one of the classes can be induced.
Demonstration.
We represent an input-label pair as $x_{1}x_{2}\cdots\to r$, where $x_{1}x_{2}\cdots$ is the input part and $r\in\{r_{+},r_{-}\}\subset\Gamma$ is an expression in $\Gamma$ serving as the verbalizer.
Chain-of-thought.
A chain-of-thought is in the same format as the training data but ends with an expression $r\in\{r_{+},r_{-}\}$, e.g., $x_{1}x_{2}\cdots\to x_{3};x_{3}\cdots x_{4}\to r_{+}$. This chain-of-thought reflects the step-by-step induction process from the inputs to the label. We show an example in Figure 1.
7.1.2 Distribution Shifts
We make experimental designs to simulate the real-world distribution shifts from training to inference:
@@ 7.1.2 Distribution Shifts
The reasoning steps are present in the training data but not in the prompts.
@@
When we choose the expressions in $\Gamma$, we assign the probability mass 45%, 45%, 5%, 5% to $r_{1},r_{2},r_{3},r_{4}$. In this way, we can inspect the effect of using less frequent verbalizers.
@@
To investigate whether the model can generalize to a new combination of formulas unseen in the training data when we generate our training data, we ensure that the expressions in $\Gamma$ are only associated with the disjunctions of two atom concepts $s_{1},s_{2}$ from a strict subset of all possible combinations $\Sigma\times\Sigma$. We then test the trained model on tasks where $v_{+}$ and $v_{-}$ are the disjunctions of the unseen combinations. We also test the models on tasks where $v_{+}$ and $v_{-}$ are the disjunctions of three atom concepts $\in\Sigma\times\Sigma\times\Sigma$.
7.1.3 Experiment Details
We use $p_{skip}=0.25$ in our experiment, generating 60,000 documents with 16 paragraphs (as described above). Among them, we use 10k for validation. We train a 6-layer Transformer (Vaswani et al., 2017) model until the loss on the validation set converges using the standard autoregressive loss. We include additional setups in §F.3.
7.1.4 Results and Discussion
@@
Table 1: The 4-shot accuracy of ICL versus chain-of-thought (CoT) using different verbalizers.
Figure 2 shows that LMs trained with Calcutec can perform in-context learning. This evidence supports our Pelican Soup framework and aligns with our theoretical analysis in §4: The ICL accuracy follows a trend $\exp(c/T)$ for some $c<0$, as our theoretical result in Eq. 1.
We further inspect the ICL performance under the distribution shifts described in §7.1.2. For the infrequent verbalizer, we observe similar performance regardless of the frequency of the verbalizers ($r_{1},r_{2}$ versus $r_{3},r_{4}$). For the unseen tasks, Figure 2 shows that the model has similar performance in tasks defined with seen and unseen combinations of atom concepts (dotted and solid lines), even generalizing to tasks defined with three latent concepts (green lines). In sum, the results show that the model can generalize well under several distributional shifts.
(a) Verbalizers = $r_{1},r_{2}$
(b) Verbalizers = $r_{3},r_{4}$
Figure 2: In-context learning accuracy with Calcutec when using different verbalizers ($r_{1},r_{2}$ or $r_{3},r_{4}$). The dotted lines represent the performance of unseen combinations described in §7.1.2. The colors represent the number of atom concepts each class ($v_{+}$ or $v_{-}$) is associated with. The main lines represent the average accuracy of 5 tasks. Lines in the lighter color represent the individual tasks.
We also experiment with 4-shot learning using chain-of-thought. Table 1 shows that the model also benefits from chain-of-thought. We conjecture that it is because chain-of-thought has a format more similar to the format for training.
7.2 Inspect Instruction-following Capability
In addition to ICL, we inspect another intriguing capability of LLMs, the capability of following unseen instructions. Our extended framework in §5 suggests that we can see following unseen instructions as doing a sentence completion task that involves atom concepts (or say “skills” in Arora and Goyal (2023)) whose composition is different from the ones involved when completing sentences in training data. We also suggest that LMs acquire the capability of following these unseen instructions by modeling the interrelations between the atom concepts involved. To show the plausibility of this, we experiment with an arithmetic task.
7.2.1 Arithmetic Task
Single-step 0 + 1 * 2 + 3 * 4 = 10 .
Multi-step 0 + 1 * 2 = 2 . 3 * 4 = 12 . 2 + 12 = 14 .
Test - seen 2 * 3 + 4 * 5 + 6 =
Test - unseen 2 + 10 + 4 * 11 * 6 =
Table 2: Training (first two rows) and test (last two rows) examples used in the arithmetic task §7.2.1.
We utilize the algebraic structure of the integers under modulo addition/multiplication to construct a language where each expression is constantly associated with a meaning and the atom concepts include 0 to 15, the symbol “+”, “*”, “=”, “.”. Each sentence in this language is in the form of an equation with additions and multiplications.
We generate LM training data in a multi-step and a single-step setup. In the multi-step setup, each example contains the steps required to compute the result, each of which involves two to five numbers (details in §G). We expect that LMs can learn the interrelation between atom concepts by modeling these steps. In the single-step setup, the final result is computed in a single step.
To simulate the scenario where completing the prompts in the test set involves a different distribution of atom concepts from that of the training examples, we ensure that the five numbers for each training example are either all from $\{0,\cdots,9\}$ or all from $\{6,\cdots 15\}$. We then evaluate the model with a seen and an unseen setups. In the seen setup, we prompt the model with 5 numbers sampled in the same way as the training set. In the unseen setup, we prompt the model to complete equations with 5 numbers where the first, third, and fifth numbers are from $\{0,\cdots,5\}$ and the other two numbers are from $\{10,\cdots,15\}$ (examples in Table 2).
7.2.2 Results and Discussion
test $\backslash$ train Multi-step Single-step
Seen 97.5 (0.3) 99.3 (0.6)
Unseen 75.3 (2.0) 49.7 (1.0)
Table 3: The average accuracy of LMs trained for the arithmetic tasks with 5 random seeds.
Table 3 shows that all LMs achieve a near-perfect accuracy in the seen setup but the LMs trained with the multi-step training set achieve significantly better accuracy in the unseen setup. This indicates that modeling the multiple steps in the training data may allow the model to learn the interrelation between the atom concepts and generalize to prompts that involve unseen compositions of atom concepts to some extent, as suggested by our extended framework in §5. This is also aligned with the success of symbolic chain-of-thought distillation (Li et al., 2023a; Hsieh et al., 2023; Shridhar et al., 2023).
task SST-2 CR MR Subj
direct 63.0 61.7 59.2 51.0
direct w/ foo/bar 51.7 53.7 52.2 57.1
pronoun 65.3 62.9 56.7 62.2